Write the quadratic equation whose roots are -2 and 5, and the leading coefficient is 2. (Use the letter x to reepresent the variable. Thank you!
Answer by oberobic(2304)

Since the roots are -2 and 5, then the obvious factors are:
(x+2)(x-5) = 0
BUT
You're told the coefficient of the x^2 term is 2.
(2x+4)(x-5) = 0
2x^2 -10x +4x -20 = 0
2x^2 -6x -20 = 0
.
Notice that you could divide through by 2.
x^2 -3x -10 = 0
(x+2)(x-5) = 0
.
Done